OpenStudy (anonymous):

Find the Lcd of X/(X-7), X/(X+7)

OpenStudy (jonnyvonny):

When you have 2 fractions, one way to get the LCD is by multiplying the denominators: \[1/4 - 1/3\rightarrow 3/12 - 4/12=-1/12\] The same principle applies: \[x/(x-7) (spacebar) x/(x+7)\rightarrow (x(x+7)(spacebar)x(x-7))/(x+7)(x-7)\] Factor that out, and you'll have a common denominator, and what the numerator is, for both equations. From then, you can add or subtract (whichever is needed).

OpenStudy (anonymous):

Thank you so much for taking the time to help me out! I factored everything out, and came up with this answer. \[\frac{ x^2+7x }{ x^2-49 }, \frac{ x^2-7x }{ x^2-49 }\] what can I do from this point?

OpenStudy (jonnyvonny):

From then, since we just wanted the GCF, we can't do anything else.

OpenStudy (jonnyvonny):

because there are no addition or subtraction signs, we can't do anything from there.